# Uncertainty-driven Adaptive Exploration

### v2

@@ 3 Experimental Setup and Results
To experimentally test our framework, we create two main instances of ADEU, using as $f(s)$ (i) the intrinsic motivation mechanism proposed in [3] and (ii) the epistemic uncertainty measuring mechanism proposed in [4]. These are the agents described in Table 1. Both agents were built on top of a TD3 [11] DRL algorithm’s original implementation. [^8] https://github.com/sfujim/TD3. As such, we refer to them as: (i) TD3+ADEU with f(s)=RNDf(s)=RND; [^9] More precisely but less concisely, “TD3+ADEU with $f(s)=r_{\text{intrinsic}}^{RND}$”. and (ii) TD3+ADEU with $f(s)=Q_{\text{std}}\left(s,\pi(s)\right)$.
@@
In addition, we consider an instance of ADEU with a non-informative $f(s)$, and create a ‘TD3 + ADEU with $f(s)=c$’ agent, where $c$ is a constant. Obviously, since $c$ is a constant, $f(s)=c$ is not an appropriate uncertainty measuring mechanism. The agent with constant $f(s)$ will serve as baseline since its distribution matches the one used in TD3 [11] and hence, they share the same exploration strategy. We use $g(f(s))=c$ for this ADEU instance. For the other ADEU instances, we use Equation 5, where $c$ has the same value as in the naive ADEU instance. That value was set to $0.2$, as in the TD3 implementation.
@@
As mentioned, our agents use TD3 as a “basis” algorithm. In principle, any DRL algorithm could have been used instead (e.g., any DRL method well-suited for continuous action spaces, if the focus is in such domains). Since ours is an exploration framework, we wanted to use a DRL algorithm that is not tied to any exploration strategy—and TD3 uses a ‘vanilla’ one that can be readily replaced with ADEU. By contrast, SAC [12] uses the entropy framework, and hence it is tied to that framework’s particular exploration strategy.
We also implement three agents to act as our main competitors, in order to test ADEU against representatives of three standard families of exploration methods: IM exploration, epistemic uncertainty-driven exploration, and (parameter-space) noise exploration.
Specifically, we test against TD3+RND, an agent that acts greedily in the environment with a reward function equal to $r_{\text{extrinsic}}(s)+r^{\text{RND}}_{\text{intrinsic}}(s)$ (cf. Equations 2 and 3). In addition, we compare against TD3+UCB, an agent that selects actions aiming to maximize Equation 4.
Finally, for interest, we also create an agent that uses Noisy Nets [10] to perturb the actor’s network and achieve more efficient exploration. We compare against this TD3+Noisy Nets agent, since Noisy Nets is a popular choice for exploring robotic domains. Note that Noisy Nets does not utilize any definition of uncertainty—it is just an “improved” TD3 agent in terms of exploration.
We conduct our experiments in standard MuJoCo domains, provided by [31]. Specifically, we used the five domains considered as the most difficult ones according to the official documentation. [^10] https://gymnasium.farama.org/environments/mujoco/ In these domains, the agent needs to explore in the continuous multidimensional action space to find the best policy. A dense reward function is used to describe these robotic tasks. However, ADEU can also be used in environments with a sparse reward function, when using appropriate uncertainty-measuring mechanisms e.g., [1, 20]. Note that we did not use [7] since their domains do not have the early termination condition.
@@
Hum. Stand. 61511; 64490 67881; 73303 80277; 97208 60878; 62227 47783; 47928 62152; 63645
Table 4: Mean and Max episodic rewards in various domains [31]. Results are averages over $15$ runs.
Table 4 presents agents’ performance in terms of mean and max episodic accumulated reward in the aforementioned domains. These mean and max values were calculated over $15$ runs (seeds), via the process described in Section 2.2. We mark in bold the best-performing algorithms in terms of mean episodic reward within each domain—an algorithm is considered to be “best-performing” if its mean episodic reward reaches at least $95\%$ of the highest mean episodic reward accumulated in the domain.
@@
As seen in Table 4, TD3+ADEU with $f(s)=RND$ or TD3+ADEU $f(s)=\lambda Q_{std}(s,\pi(s))$ outperform the others in every domain. In addition, in the three most difficult domains (i.e., Ant, Humanoid Standup, and Humanoid) both of those ADEU instances outperform all other agents, achieving the first and the second best performance. It is also noteworthy that an instance of ADEU equipped with a specific uncertainty-measuring mechanism $f(s)$, typically outperforms the original exploration method that employs the corresponding uncertainty-measuring mechanism.
In addition, our results show that ADEU combined with $\lambda Q_{\text{std}}(s,\pi(s))$ outperforms most, if not all, of its opponents in the highly challenging Ant, Humanoid, and Humanoid Standup environments. That indicates that this mechanism tracks uncertainty more efficiently in larger environments.
@@
As seen in our experiments, ADEU can incorporate any uncertainty-measuring mechanism of choice. Of course, as also discussed above, not all uncertainty-measuring mechanisms are created equal. Notice, e.g., that the naive instance of TD3+ADEU, TD3+ADEU with $f(s)=c$, is outperformed by at least one other ADEU instance in each domain. This indicates that the efficiency of ADEU depends on the efficiency of the uncertainty mechanism. Seen otherwise, the choice of an appropriate uncertainty-measuring mechanism for use in the ADEU framework, promises the improvement of performance in the domain of interest.
@@ 4 Related Work
Intrinsic Motivation: Several IM approaches have been tested in robotic domains. For instance, Lobel et al. [17] employ a Coin Flip Network to design an efficient intrinsic motivation function that can be utilized as an uncertainty mechanism in ADEU. They test their method in a sparse reward variance of the Fetch domain, which is more comparable to deep exploration settings. On the other hand, [18] adds a random reward to the sparse reward function of the environment, to guide the agent to explore different parts of the environment. This can be used for testing our framework in sparse reward function domains.
Epistemic Uncertainty: As mentioned earlier, UCB [4, 13] is commonly used in continuous action-space robotic domains. Inspired by these works, we also created an instance of ADEU using this ensemble mechanism to track uncertainty (the ADEU instance with $f(s)=\lambda Q_{\text{std}}(s,\pi(s))$ shown in Section 3) and test it against a UCB implementation built on top of a TD3 agent. Similarly, [5] measures uncertainty as the variance of the different $Q$ representations. In the case of high uncertainty, the ensemble agent “asks” a mentor for guidance.
Adaptive exploration: Tokic et., al [29], uses the Temporal-Difference Error to change the $\epsilon$ parameter of the $\epsilon$-greedy in each state. Hence, an agent is more likely to explore when uncertainty about the $Q$-function is high. Pislar et., al [23] is the first work that actively attempts to create a modern framework for intra-episodic exploration. They argue that an agent should switch from exploitation to exploration when an external signal arrives. They propose numerous “signals” for changing modes, either heuristic or random ones. As noted earlier, these signals can be used as uncertainty mechanisms in ADEU. Similarly to [23], ez-greedy [6] suggests following directed behavior with random—sampled from a heavy-tailed distribution—length. This directed behavior, guides the agent to repeat a single action numerous times. Thus, it achieved excellent results in settings which require replicating the same action numerous times, such as DeepSea [22] and Chain Walk [21]. On a similar note, [16] proposes an adaptive exploration variant of PPO [24], where the entropy term in the loss function is scaled by a heuristic based on the agent’s recently obtained rewards. Even though ADEU does not directly manipulate the loss, it can use this heuristic as $f(s)$. Finally, most goal-based exploration works [8, 9, 30] can be considered as ‘adaptive exploration’ ones since they replicate the known trajectory with no exploration, and then use some exploration to extend it.
@@ 5 Conclusions and Future Work
In this paper, we put forward a generic adaptive exploration framework, that uses a generic uncertainty-based action selection mechanism to decide in a principled manner when to alternate between exploration and exploitation and vice versa. The generality of that mechanism allows (i) our framework to effectively incorporate any uncertainty-measuring mechanism of choice; and (ii) existing adaptive exploration techniques to be viewed as special cases of our framework. In problems requiring adaptive exploration, ADEU is a promising and easily deployable solution. Additionally, it allows the user to select an existing uncertainty measurement mechanism or to define a heuristic one tailored to its problem. Our experiments verified that ADEU instances outperform other exploration strategies in standard difficult robotic testbeds.
Ongoing and future work includes performing more tests and extending ADEU in various directions. To begin, we intend to perform tests to further verify the ability of ADEU to exploit background knowledge; and its ability to recover from being fed with sub-optimal policies and erroneous background information (e.g., via the use of the rollout episodes or via designing mechanisms that account for a policy’s value). Finally, we intend to design mechanisms to appropriately set the third and fourth moments of $D$, to guide exploration to particular directions—a useful feature for safe exploration.

### v3

@@ 3 Experimental Evaluation
To experimentally test our framework in earnest, we create two main instances of ADEU, using as $f(s)$: (i) the intrinsic motivation mechanism proposed in burda2018rnd; and (ii) the epistemic uncertainty measuring mechanism proposed in Chen2017ucb. These are the agents described in Table 1. Both agents were built on top of a TD3 [fujimoto2018TD3] DRL algorithm’s original implementation. [^8] https://github.com/sfujim/TD3. As such, we refer to them as: (i) TD3+ADEU with f(s)=RNDf(s)=RND; [^9] More precisely but less concisely, “TD3+ADEU with $f(s)=r_{\text{intrinsic}}^{RND}$”. Reward function choices and implementation details for all agents are in the Appendix. and (ii) TD3+ADEU with $f(s)=\lambda Q_{\text{std}}\left(s,\pi(s)\right)$.
@@
In addition, we consider an instance of ADEU with an uninformative $f(s)$, and create a “TD3 + ADEU with f(s)=cf(s)=c” agent, where $c$ is a constant. Obviously, $f(s)=c$ is not an appropriate uncertainty measuring mechanism. The agent with constant $f(s)$ will serve as baseline, since its distribution matches the one used in TD3 [fujimoto2018TD3] and hence, they share the same exploration strategy. We use $g(f(s))=c$ for this ADEU instance.
For the other ADEU instances, we use Eq. 12, with $c=0.2$ as in TD3 implementation. (More details regarding the choice of $g(\cdot)$ are in the Appendix.)
@@
Table 5: Mean and Max episodic rewards in various domains [towers2024gymnasium]. Results are averages over $15$ runs.
@@
Hum. Stand. 61511; 64490 73895; 79978 72978; 80139 54928; 56527 55211; 56057 62152; 63645
Our agents use TD3 as a “basis” algorithm. In principle, any DRL algorithm could have been used instead. Since ours is an exploration framework, we used a DRL algorithm not tied to any exploration strategy—and TD3 uses a ‘vanilla’ one that can be readily replaced with ADEU. By contrast, SAC [Haarnoja2018SAC] uses the entropy framework, and hence, it is tied to that framework’s particular exploration strategy.
We also implement three agents to act as our main competitors, to test ADEU against representatives of three standard families of exploration methods: IM exploration, epistemic uncertainty-driven exploration, and (parameter-space) noise exploration. Specifically, we test against TD3+RND, an agent that acts greedily in the environment with a reward function equal to $r_{\text{extrinsic}}(s)+r^{\text{RND}}_{\text{intrinsic}}(s)$ (cf. Equations 2 and 3). In addition, we compare against TD3+UCB, an agent that selects actions aiming to maximize Equation 4.
Finally, for interest, we also create an agent that uses Noisy Nets [fortunato2018noisy] to perturb the actor’s network and achieve more efficient exploration. We compare against this TD3+Noisy Nets agent, since Noisy Nets is a popular choice for exploring robotic domains. Note that Noisy Nets does not utilize any definition of uncertainty—it is just an “improved” TD3 agent in terms of exploration.
We conduct our experiments in standard MuJoCo domains [towers2024gymnasium]. Specifically, we used the five domains considered the most difficult according to the official documentation. [^10] https://gymnasium.farama.org/environments/mujoco/ . We did not use [gymnasium_robotics2023github] since its domains do not have the “early termination” condition. Such domains do not satisfy Definition 2.1. However, the ADEU agents still achieve (as later described) superior performance against their opponents. In these domains, the agent should explore the continuous multidimensional action space to find the best policy. A dense reward function is used to describe these robotic tasks. However, ADEU can also be used in environments with a sparse reward function, when using appropriate uncertainty-measuring mechanisms e.g., [Badia2020NeverGiveUp, nikolov2018informationdirected].
Table 5 presents agents’ performance in terms of mean and max episodic accumulated reward in the aforementioned domains. These mean and maximum values were calculated over $15$ runs (seeds), via the process described in the motivation example. We present both maximum and mean reward to track both the best performance of an algorithm and an indication of the time required to reach convergence. We mark in bold the best-performing algorithms in terms of mean episodic reward within each domain—an algorithm is considered to be “best-performing” if its mean episodic reward reaches at least $90\%$ of the highest mean episodic reward accumulated in the domain. The (standard) figures where the cumulative episodic evaluation reward is plotted can be found in the Appendix.
@@
As seen in Table 5, TD3+ADEU with $f(s)=RND$ or TD3+ADEU $f(s)=\lambda Q_{std}(s,\pi(s))$ outperform the others in every domain. In addition, in the three most difficult domains (i.e., Ant, Humanoid Standup, and Humanoid) both of those ADEU instances outperform all other agents, achieving the first and the second best performance. It is also noteworthy that an instance of ADEU equipped with a specific uncertainty-measuring mechanism $f(s)$ typically outperforms the original exploration method that employs the corresponding uncertainty-measuring mechanism.
In addition, our results show that ADEU with $\lambda Q_{\text{std}}(s,\pi(s))$ outperforms most of its opponents in the three aforementioned highly challenging environments. This indicates this mechanism tracks uncertainty more efficiently in larger environments.
@@
As seen in our experiments, ADEU can incorporate any uncertainty-measuring mechanism of choice. Of course, as also discussed above, not all uncertainty-measuring mechanisms are created equal. Notice, e.g., that the naive instance of TD3+ADEU with $f(s)=c$ is outperformed by at least one other ADEU instance in each domain. This indicates that ADEU efficiency depends on the efficiency of the uncertainty mechanism. Seen otherwise, the choice of an appropriate uncertainty-measuring mechanism for use in ADEU, promises the improvement of performance in the domain of interest.
Results in the DeepSea domain
[…]
Figure 4: Timesteps required for each algorithm to solve the DeepSea domain for $N\times N$ grid. The lower the curve, the better the performance of the algorithm.
[…]
From the conducted experiments and the corresponding results, it is clear that the ADEU framework, paired with efficient uncertainty-measuring mechanisms, can be utilized for both demanding control tasks and domains requiring deep exploration.
@@ 4 Related Work
We now briefly discuss related work regarding IM, epistemic uncertainty, or adaptive exploration.
Intrinsic Motivation: Several IM approaches have been tested in robotic domains.
For instance, [Lobel2023FlippingCoins] design an efficient intrinsic motivation function that can be utilized as an uncertainty mechanism in ADEU. They test their method in a sparse reward variance of the Fetch domain; an easier domain than standard robotics. Then, [mahankali2024random] adds a random reward to the sparse reward function, to guide the agent to explore different parts of the environment. This can be used by ADEU in extremely sparse reward function domains.
Epistemic Uncertainty: As mentioned earlier, UCB [Chen2017ucb, lee2021sunrise] is commonly used in continuous action-space robotic domains.
We also created an instance of ADEU using this ensemble mechanism to track uncertainty (the ADEU instance with $f(s)=\lambda Q_{\text{std}}(s,\pi(s))$ in Sec. 3) and test it against a UCB implementation built on top of a TD3 agent. da2020uncertainty “asks” a mentor for guidance when the uncertainty is high; uncertainty was measured in a similar way.
Active Exploration: Several works [russo2023active1, hazan2019active2, hong2020active4, nikolov2018informationdirected] have proposed methods to efficiently explore the environment to derive an $\epsilon$-optimal policy, while also obtaining a low sample complexity. Some methods [russo2023active1, hazan2019active2], where tested in tabular or “discretized” environments. Hence, these works tackle a distinct exploration problem, and there is no straightforward comparison with ADEU. Other works [hong2020active4, nikolov2018informationdirected] incorporate these ideas in the classic DRL testbeds either by proposing mechanisms to motivate efficient exploration [nikolov2018informationdirected] or by changing the reward function that the agent aims to maximize [hong2020active4]. Both mechanisms can be used in the ADEU framework.
[…]
@@ 5 Conclusions and Future Work
In this paper, we put forward a generic adaptive exploration framework that uses a generic uncertainty-based action selection mechanism to decide in a principled manner when to alternate between exploration and exploitation. The generality of that mechanism allows (i) our framework to effectively incorporate any uncertainty-measuring mechanism of choice; and (ii) existing adaptive exploration techniques to be viewed as special cases of our framework. In problems requiring adaptive exploration, ADEU is a promising and easily deployable solution. Additionally, it allows the user to select an existing uncertainty measurement mechanism or to define a heuristic one tailored to its problem.Our experiments verified the superiority of ADEU in difficult robotic testbeds.
Ongoing and future work includes extending ADEU in various directions. To begin, we intend to perform tests to further verify ADEU’s ability to exploit background knowledge; and its ability to recover from being fed with incorrect background information or to detect and escape clear suboptimal trajectories by optimizing the rollout procedure. In addition, to make the learning procedure more robust, we will modify the learning process by weighting the learned experience before backpropagating it in the network, similar to [lee2021sunrise]. As mentioned earlier, we did not modify the learning procedure in this work, since ADEU does not exhibit instabilities, and we also aim to keep this work “exploration-only”. Moreover, we are actively working to extend ADEU to guarantee safe exploration [wachi2023safe, ijcai2022p3o]. In ongoing work, we are calculating the upper bound of allowed exploration in neighboring safe states, so that ADEU can guarantee safe exploration. We also aim to extend ADEU to multiagent environments. By leveraging agents’ independent exploration, ADEU agents can adaptively explore the environment. Furthermore, by allowing a centralized agent to define $f_{i}(s)$ for each agent $i$, we can reduce the non-stationarity issues that usually arise in multiagent RL. Finally, we intend to design mechanisms to appropriately set the third and fourth moments of $D$, to guide exploration to particular directions.
@@
Appendix A Technical details regarding the “modified Frozen Lake” motivating example
For our motivation example, we create an environment similar to Figure 5, but we increase the dimensions of the grid to $1500\times 1500$. As mentioned in the main text, we test $4$ different agents in the modified (as described in the main paper) Frozen Lake environment: [^11] We note that all agents were also tested in the classic Frozen Lake, for fine-tuning debugging reasons and for fine-tuning their parameters. Hence, we do not present these results here. (i) the classic Q-learning agent paired with $\epsilon$-greedy; (ii) an IM agent built on top of the Q-learning agent; (iii) a UCB agent consisting of $5$ Q-learning agents; and (iv) an ADEU agent built on top of the Q-learning agent, with an $f(s)$ as described below.
Figure 5: Environment used in our motivation example. In this particular figure, the dimensions of the grid are $10\times 10$. In our example, we use a $1500\times 1500$ grid.
As discussed in the main text, in our motivation example, we used a dense reward function that rewards the agent when it reaches a state closer to the goal. In more detail, we reward the agent every time it performs a ‘forward’ step and reaches a valid, non-terminal state $s$ according to Equation 8:
$r(s)=1500\cdot\frac{s}{\sum_{i=0}^{i=1500^{2}}i}$ (8)
where $s$ is the id of the current state (i.e., a number between 0 and $1500^{2}-1$); and $\sum_{i=0}^{i=1500^{2}}i$ serves as a normalizer.
In the case of a “hole”, the agent receives a penalty (=10)(=-10), and the episode is terminated; while when the agent reaches the goal terminal state, it receives a high reward—equal to $10^{5}$ in our case.
The theoretical maximum reward in this environment when $N=1500$ is calculated as the sum of the rewards in the states of the sole trajectory i.e., states with: idid 0 and idsids between $1$, and (N2)N+2(N-2)N+2 with step $N$ and idsids between (N2)N+2(N-2)N+2 and (N2)N+N(N-2)N+N and state $(N^{2})-1$ ($2999$ in total). The evaluation reward for those states is $104,489$ (unsmoothed—$4,489$ for the trajectory + $10^{5}$ for finding the treasure) or $104,487$ (smoothed), equal to the one obtained by ADEU (see Table 3).
As an intrinsic reward mechanism for the IM agent, we use the visitation frequency, calculated as $\frac{\beta}{\sqrt{n(s)}}$, where is the exact visitation count $n(s)$. This approach is similar to [tang2017exploration], but without the need to approximate the visitation frequency, as in this tabular scenario we can in fact keep track of the times that an agent has visited a state $s$. As such, the reward function used for the IM agent is :
$r(s)=r_{\text{extrinsic}}(s)+r_{\text{intrinsic}}(s)=r_{\text{extrinsic}}(s)+\frac{\beta}{\sqrt{n(s)}}$ (9)
where $\beta$ is a constant number between 0 and $1$, and $n(s)$ is the aforementioned visitation frequency.
For the ADEU agent used in the motivation example, we select as $D$ a categorical distribution with $K$ categories, for which the probability of sampling-out an action $a$ (out of $K=4$ possible actions) at a specific state $s$ is determined by:
$p(s,a)=sigmoid\left(\beta\sqrt{n(s)}\right)\cdot q(s,a)$ (10)
where $q(s,a)$ is the $Q$-values for a state $s$, $\beta$ is a constant number between 0 and $1$, and $n(s)$ is the aforementioned visitation count. We note that $p(s,a)$ will be normalized before action selection. Note that, for the categorical distribution, if the agent’s uncertainty is high at $s$, all $p(s,a)$ should have similar values, close to $1/K$; while if the agent’s uncertainty is low at $s$ the $a$ with the highest $p(s,a)$ will be regularly selected. [^12] The variance of observing the $X_{i}$ random variable corresponding to selecting $a$ is quite small in this case—specifically, equal to p(s,a)(1p(s,a))p(s,a)(1-p(s,a)).
Now, if the agent has visited a state $s$ numerous times, and hence $n(s)\rightarrow\infty$, it will select actions with probability $p(s,a)=q(s,a)$. This is almost equivalent to the Boltzmann exploration with $T=1$. Consider now the case $n(s)=0$. The ADEU agent should uniformly sample an action in this case, to achieve full exploration. However, if $n(s)=0$, then if the classic sigmoid function is used, it holds that $p(s,a)$ is bounded by $0.5q(s,a)$ (since $0.5$ is the lower bound of the classic sigmoid for positive input).
To mitigate this issue, we employ a different sigmoid function (we henceforth refer to this as “shifted sigmoid”) that shifts the input before applying the classic sigmoid. We selected this “shifted” sigmoid since it has a lower than $0.5$ lower bound even when $x$ is positive. This will result to pre-normalized $p(s,a)$ values that are close to 0 when $n(s)=0$; and, thus, following normalization to $p(s,a)$ values that are close to $1/K$ each for a particular $s$. Hence, it allows for more exploration when the visitation count of a state is low.
That is, for our ADEU agent in this motivation example, we sample actions at each state according to:
$p(s,a)=shifted\_sigmoid\left(\beta\sqrt{n(s)}\right)\cdot q(s,a)$ (11)
Figure 6 shows the two sigmoid functions.
Figure 6: The “shifted” sigmoid function compared to the classic sigmoid.
Appendix B Ablation Study on the Normalizer Function
Eq. 12 shows the $g(\cdot)$ used in our main experiments.
$g(f(s))=sigmoid(f(s))\cdot c$ (12)
Since $c=0.2$, $g(x)$ is bounded between $[0.1,0.2]$, given that $x=f(s)$ is always positive. We also explored the possibility of using a custom sigmoid function, $\hat{g}(x)$:
$\hat{g}(x)=(1-\exp{(-x)})\cdot c$ (13)
where $x=f(s)$ and again $c=0.2$.
Figure 7 shows the (positive input part of the) aforementioned functions: the custom sigmoid $\hat{g}(x)$ will output a 0 value in case of a low positive input value—i.e., when that agent has no uncertainty ($f(s)=0$), the variance of the distribution would be set to 0.
Figure 7: The two sigmoids tested in our main experiments
While this approach has desirable properties, it may lead to unintended consequences. Specifically, when uncertainty is low, the agent may repeatedly generate identical or highly similar experiences. This lack of diversity in the training data can increase the risk of catastrophic forgetting, since the agent produces correlated experience [Bruin2015catastrophic_forgetting_2]—at least for the trusted policy.
This is why in our experiments we resolved to not using the custom sigmoid function, but use the function of Equation 12 as our $g(\cdot)$, as previously indicated. Since the primary objective of our work in this paper is to address in a principled manner the problem of adaptive exploration, we were content with selecting a suitable form for $g(\cdot)$, i.e., that of Equation 12; and did not attempt to move the focus away from exploration itself in order to optimize performance. Regardless, in future work, we plan to re-visit the $\hat{g}$ approach and attempt to avoid the catastrophic forgetting problem as follows.
[…]
Appendix C Implementation details on RND’s networks
As discussed in the main text, the novelty detection mechanism used in RND for providing the intrinsic reward consists of two networks: (i) the Target; and (ii) the Predictor. Both networks use a state $s$ as input. The output of the Target network is a constant but randomly selected value, while the Predictor aims to predict it. Eventually, for the non-novel states, Predictor will converge to the Target’s output. The convergence speed is highly affected by the hyperparametrs of the two networks. Instead of modifying these numerous parameters (e.g., learning rate of the Predictor, number of layers of each network, type of the optimizer etc) we have introduced a parameter that scales the output of both networks, termed $U$. That is, the final uncertainty-measuring mechanism used in the corresponding ADEU instance is the following:
$f_{\text{RND}}(s)=r^{\text{RND}}_{\text{initrinsic}}=\left\lVert U\cdot f_{\text{predictor}}(s)-U\cdot f_{\text{target}}(s)\right\rVert_{2}^{2}$ (14)
We observe the behavior of the ADEU instance for numerous values of $U$, and we selected $U=187.5$. This value was set for all the experiments. [^13] In addition, we create a mechanism able to dynamically fine-tune $U$ throughout the training process by decreasing it if $f_{\text{RND}}(s)$ is high and increasing it if it is too low. We observed that this heuristic modification was used by ADEU only in the later stages of training (approximately after $85\%$ of the training). We note that this modification (employing $U$) was not used in the TD3+RND agent since this would have affected the robustness of the agent.
Appendix D Hyperparameters and Hardware details
Table 6 lists all hyperparameters used in our experiments. To fine-tune the hyperparameter $\lambda$ for both TD3+UCB and ADEU with $f(s)=\lambda Q_{\text{std}}(s,\pi(s))$, we conducted experiments in the Hopper domain. We selected $\lambda=10$ for TD3+UCB, as this value led to the best performance for TD3+UCB in this domain. For ADEU, we chose $\lambda=1$ for the same reason. We note however that even with $\lambda=10$, ADEU with $f(s)=\lambda Q_{\text{std}}(s,\pi(s))$ consistently outperformed TD3+UCB (with $\lambda=10$) in this environment.
Furthermore, we compared ADEU using $f(s)=\lambda Q_{\text{std}}(s,\pi(s))$ and $f(s)=f_{\text{RND}}$ in the Hopper domain. Based on our results in this domain, we set $\rho=0.1$ for the former, and $\rho=0.3$ for the latter. We also observe improved behaviour in experiments in more difficult domains (such as the Humanoid etc) when $\rho=0$ and $f(s)=\lambda Q_{\text{std}}(s,\pi(s))$. We did not conduct experiments with $\rho=0$ and $f(s)=f_{\text{RND}}$ in the difficult domains. However, we only report the results for $\rho=0.1$ in the case of $f(s)=\lambda Q_{\text{std}}(s,\pi(s))$. In future work, we plan to further investigate the behaviour of the agent for different values of $\rho$.
Table 6: Parameters used in our experiments
Hyperparameter Value
Discount ($\gamma$) 0.99
Soft update rate ($\tau$) 0.005
Policy noise ($c$) 0.2
Noise clip 0.5
Policy update frequency 2
Learning rate 0.00005
Batch size 256
Buffer size $10^{5}$
$U$ 187.5
$\lambda$ [1, 10]
$\rho$ [0.1, 0.3]
$\beta$ (for UCB’s binomial) 1
All of our code was implemented in Python using the $3.11$ version. In addition, we use the $2.3.1$ version of pyTorch [paszke2019pytorch] and the $1.26.1$ version of numpy [harris2020numpy]. We also use the $1.1$ version of gymnasium [towers2024gymnasium], and we ran our experiments using the latest ($v5$) version of the domains.
Appendix E Extended results
Figure 8 shows the performance of each exploration strategy/framework in the tested domains. These results were used to calculate Table 4 of the main text.
Figure 8: The cumulative episodic rewards yielded by each exploration strategy/framework in the tested domains. Results are averaged over $15$ seeds. Shaded areas correspond to the standard deviation.
Appendix F Recreating ez-greedy as an ADEU instance
As already discussed, a Distribution $D$, a policy $\pi(s)$, and an uncertainty-measuring mechanism $f$ are required, to define any instance of an ADEU agent. To emulate dabney2021ez_greedy, we will define the following parameters:
Table 7: Emulating $\epsilon z$-greedy
Parameter Selection
Distribution $D$ Bernoulli distribution.
Policy $\pi(s)$ Equation 15
Uncertainty-Measuring Mechanism Algorithm 1
Equation 15 describes the stochastic policy of the ADEU agent that emulates the $\epsilon z$-greedy algorithm. In more detail, the agent utilizes a Bernoulli distribution that selects the “exploitation” action with probability $1-f(s,n,\cdot)$, or follows the given option for $n$ timesteps. Intuitively, ADEU selects an action that promotes exploration when the uncertainty $f(s,n,\cdot)$ is high.
$\pi(s)=\left\{\begin{matrix}\arg\max_{i}Q(s,a_{i}),\quad 1-f(s,n,\cdot)\\ \omega\sim U(A),\quad f(s,n,\cdot)\end{matrix}\right$. (15)
where $f(s,n,\cdot)$ is the uncertainty-measuring mechanism described in Algorithm 1, $Q(s,a_{i})$ is the discrete action space Q-Function for each action $i$ derived from any DRL algorithm of choice, and $U(A)$ is the mechanism that calculates the options as described in [dabney2021ez_greedy].
[…]
1:Inputs: $\epsilon$ (float): probability to begin an option (see [dabney2021ez_greedy]), $n$ (integer $\geq 0$): steps exploring using an option, explored (boolean): indicates whether the agent explored or not during the previous step.
2: Output: uncertainty, $n$ (integer $\geq 0$) //$n$ is returned for internal/“local” use
3:if $n==0$ and not explored then
4:  // Begin an option with probability $\epsilon$
5:  return $\epsilon,0$
6:end if
7:if $n==0$ and explored then
8:  // The agent has already utilized the option just one time (i.e., in the previous timestep the agent sampled the $U(A)$ action in Equation 15)
9:  sample $n$
10:  return 1,n11,n-1
11:end if
12:if $n>0$ then
13:  // Keep exploring using the option
14:  return 1,n11,n-1
15:end if
[…]
To be consistent with the previous notation, in that case $f()$ is called with n=n=previous returned $n$, $\epsilon=\rho$, and explored = True if the agent sampled $U(A)$, otherwise it is set to False. For the rest “implementation details” please refer to the corresponding Section and dabney2021ez_greedy.
[…]
Theorem 2.
[…]
$pr=1-(1-p)^{E}$
$\text{where }p=\left(\nicefrac{{1}}{{N+1}}\right)\left(1-\frac{\sum_{i=1}^{N}i^{-s}}{\zeta(s)}\right)\text{ and, }$
$s=\mu$
Proof.
[…]
$p=\nicefrac{{\epsilon}}{{2}}Pr(n\geq N)=$
$=\nicefrac{{\epsilon}}{{2}}Pr(n+1)>N=$
$=\nicefrac{{\epsilon}}{{2}}\left(1-Pr(n+1\leq N)\right)\overset{\underset{\epsilon=\nicefrac{{1}}{{N+1}}}{}}{=}$
$=\left(\nicefrac{{1}}{{2(N+1)}}\right)\left(1-\frac{\sum_{i=1}^{N}i^{-s}}{\zeta(s)}\right)$
since (from the definition of the Zeta distribution)
$Pr(n+1\leq N)=\frac{H_{n,s}}{\zeta(s)}$
$\text{where }H_{n,s}=\sum_{i=1}^{N}i^{-s}$
$\text{and }s=\mu$
Previously, we calculated the probability of sampling the ‘go-left’ option with a length greater than the size of the grid, in a one-trial experiment. However, this experiment will be conducted $E$ independent times. Thus, to calculate the probability of selecting the optimal trajectory at least once, we will construct a Binomial distribution $B(E,p)$. Then, we will calculate the Cumulative distribution function for $Pr(X\geq k)$ for $k=1$.
$Pr(X\geq 1)=1-P(X\leq 0)=$
$=1-\sum_{i=0}^{\lfloor 0\rfloor}\binom{E}{i}p^{i}(1-p)^{E}-i=$
$=1-\binom{1}{0}p^{0}(1-p)^{E}=$
$=$ $1-(1-p)^{E}$
As expected, the probability is proportional to the training time $E$. Finally, by combining these, we derive the lower bound of the probability of selecting the optimal trajectory at least once:
$pr=1-\left(1-\left(\nicefrac{{1}}{{2(N+1)}}\right)\left(1-\frac{\sum_{i=1}^{N}i^{-\mu}}{\zeta(\mu)}\right)\right)^{E}$
We should highlight that Theorem 2 calculates the lower bound of finding the treasure. This bound is calculated assuming that the agent’s policy (derived by $\arg\max_{a}Q\{(s,a)\}$) always guides the agent to the wrong trajectory. In this case, $f(s,n,\cdot)$ forces the agent to explore the correct trajectory. Hence, we can easily assume that the probability of learning the correct policy increases as the agent explores the correct path.
[…]